What is the resistance (impedance) of a 3 x 10^-6 farad capacitor powered by an AC power supply with a frequency of 60 Hz?


1
Expert's answer
2021-02-07T19:09:33-0500

"X_C=\frac{1}{2 \pi f C }"


"f=60\: Hz"


"C=3 \times 10^{-6}\: F"

"\therefore X_C=\frac{1}{2*\frac{22}{7} *60*3*10^{-6}} = 883.83\:\Omega"